OpenStudy (anonymous):

what is the value of h(z)+4z^2 /5z when is z=4

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Demonx341

OpenStudy (anonymous):

How u make a question i had made one and im new here and i really donr know how to make a qustion

OpenStudy (dayakar):

\[h(z) =\frac{ 4^z{2} }{ 5z }\]

OpenStudy (dayakar):

r u typed + instead of =

OpenStudy (dayakar):

\[h(4) = \frac{ 4*4^{2} }{ 5*4 }\] since z=4

OpenStudy (dayakar):

4 is cancelled h(4)= 16/5

OpenStudy (dayakar):

get it

OpenStudy (anonymous):

so the answer ia 16/5

OpenStudy (anonymous):

@dayakar